OpenStudy (anonymous):

(6)Add and simplify if possible: 3 5 ______ + _________ x-5 x^2-25 a.) 8 _____________ (x+5)(x-5) b.) 3x+20 ______________ (x+5)(x-5) c.) 8x+5 ____________ (x+5)(x-5)

OpenStudy (lalaly):

did u try to do it ? ..

OpenStudy (anonymous):

I tried, I dont know how to solve it.

OpenStudy (anonymous):

Can you factor: x^2-25

OpenStudy (anonymous):

(x - 5)(x + 5)

OpenStudy (anonymous):

Yep :) So what's common denominator?

OpenStudy (anonymous):

x-5, right?

OpenStudy (anonymous):

I mean what should you multiply with the first fraction to get (x - 5)(x + 5)

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

Look at the first denominator, it already had ....

OpenStudy (anonymous):

x-5. I said that though

OpenStudy (anonymous):

what should you MULTIPLY with (x-5) --> COMMON DENOM (x - 5)(x + 5)

OpenStudy (anonymous):

=> MULTIPLY both upper and lower of the first fraction with (x + 5), will you!

OpenStudy (anonymous):

\[\frac{3}{x-5}+\frac{5}{x^2-25}\\\\=\frac{3(x+5)}{(x-5)(x+5)}+\frac{5}{(x-5)(x+5)}\\\\=\frac{3x+20}{(x-5)(x+5)} \]

OpenStudy (anonymous):

@qwertyfreak perform multiply in the first fraction of the second row as in @Luis_Rivera post! Learn it for your own benefit in the long run!